# Living at the Edge: A Critical Look at the Cosmological Neutrino Mass Bound

### v1

@@ Abstract
Cosmological neutrino mass bounds are becoming increasingly stringent. The latest limit within $\Lambda$CDM from Planck 2018+ACT lensing+DESI is $\sum m_{\nu}<0.072\,{\rm eV}$ at 95% CL, very close to the minimum possible sum of neutrino masses ($\sum m_{\nu}>0.06\,{\rm eV}$), hinting at vanishing or even “negative” cosmological neutrino masses. In this context, it is urgent to carefully evaluate the origin of these cosmological constraints. In this paper, we investigate the robustness of these results in three ways: i) we check the role of potential anomalies in Planck CMB and DESI BAO data; ii) we compare the results for frequentist and Bayesian techniques, as very close to physical boundaries subtleties in the derivation and interpretation of constraints can arise; iii) we investigate how deviations from $\Lambda$CDM, potentially alleviating these anomalies, can alter the constraints. From a profile likelihood analysis, we derive constraints in agreement at the $\sim 10\%$ level with Bayesian posteriors. We find that the weak preference for negative neutrino masses is mostly present for Planck 18 data, affected by the well-known ‘lensing anomaly’. It disappears when the new Planck 2020 HiLLiPoP is used, leading to significantly weaker constraints. Additionally, the pull towards negative masses in DESI data stems from the $z=0.7$ bin, which is in $\sim 3\sigma$ tension with Planck expectations. Without these outliers, and in combination with HiLLiPoP, the bound relaxes to $\sum m_{\nu}<0.11\,{\rm eV}$ at 95% CL. The recent preference for dynamical dark energy alleviates this tension and further weakens the bound. As we are at the dawn of a neutrino mass discovery from cosmology, it will be very exciting to see if this trend is confirmed by future data.
@@ I Introduction: Neutrino Mass bounds as of mid 2024
Clearly, the current cosmological limit is very close to the minimal value in normal ordering and already disfavours to some degree the inverted one. Importantly, the limit in Eq. (1) is so close to the minimum physical boundary that statistical statements about the neutrino mass need to be taken with care. However, cosmological limits are typically derived within a Bayesian framework, and as such will depend upon the priors used. While the dependence may be weak when the likelihood largely dominates over the prior, and far away from physical boundaries, it has already been established that neutrino masses are strongly sensitive to the choice of prior Simpson et al. (2017); Gariazzo et al. (2018, 2023), even with DESI Adame et al. (2024a). Moreover, and perhaps even more surprising, there is still no hint of a non-zero mass in the posterior probability density, and in fact, it has been argued that cosmological data may favor a negative effective neutrino mass Craig et al. (2024); Green and Meyers (2024); Elbers et al. (2024). While this could be the result of a statistical fluctuation or a systematic effect, it could potentially be the indication of new phenomena in cosmology with groundbreaking implications.
@@
The rest of our study is structured as follows: first, in Section II we briefly review the main cosmological implications of massive neutrinos. We will highlight the crucial role of Planck data as well as why BAO data can significantly tighten the constraints on the neutrino mass. In Section III we outline the various data sets that we will use, and describe the statistical procedure (both frequentist and Bayesian) that we use to analyse the data and derive bounds. In Section IV we present our main results focusing first on analyses with Planck data only, exploring in detail the role of the lensing anomaly; we then perform comprehensive analysis of Planck and BAO data, comparing in particular DESI and SDSS; and finally, check the impact of considering extensions to $\Lambda$CDM. In Section V we specifically address the potential cosmological preference for a negative neutrino mass. Finally, in Section VI we draw our conclusions.
@@ II Cosmological Impact of the neutrino mass
The implications of the neutrino mass in cosmology have been explored and discussed in depth and length, see Lesgourgues et al. (2013); Dolgov (2002); Lesgourgues and Pastor (2006); Hannestad (2010); Wong (2011); Lattanzi and Gerbino (2018) for reviews. In essence, the cosmological implications of neutrino masses can be understood as follows: 1) Neutrinos are always a relevant component of the energy density of the Universe and therefore contribute to its expansion rate, $H\propto\sqrt{\rho}$. 2) Neutrinos were propagating ultrarelativistically until the Universe cooled down to $T_{\nu}\simeq m_{\nu}/3$ which occurs at $z_{\nu}^{\rm NR}\simeq 190\,m_{\nu}/(0.1\,{\rm eV})$. 3) As a result of their ultrarelativistic speeds, neutrinos are not able to cluster on scales smaller than $L\simeq 2\,{\rm Gpc}\sqrt{0.1/m_{\nu}}$ today. In consequence, and for massive neutrinos which become non-relativistic after recombination $m_{\nu}\lesssim 0.6\,{\rm eV}$, one of their main cosmological implications is to suppress the amount of structure formation on scales smaller than $L\simeq 2\,{\rm Gpc}\sqrt{0.1/m_{\nu}}$. This suppression is of course strongly dependent upon the energy density neutrinos represent in the Universe which is directly proportional to the neutrino masses as given in Eq. (4). This discussion clearly highlights that the best way to search for the neutrino mass is arguably by directly observing probes of the Large Scale Structure in the Universe. This is one of the primary goals of the ongoing DESI and Euclid surveys (see Font-Ribera et al. (2014); Amendola et al. (2013); Aghamousa et al. (2016)). Nevertheless, as of today the bound on the neutrino mass is dominated by Planck with the aid of BAO and Supernova measurements to break relevant parameter degeneracies.
@@
Importantly, the CamSpec and HiLLiPoP likelihood have been recently updated since the 2018 analysis, in light of a new set of maps produced by the Planck collaboration called NPIPE. The NPIPE maps exploit a number of improvements in the processing of time ordered data to allow for an increase in the signal-to-noise ratio at small scales. They also allow to use a larger sky fraction, and incorporate a better handling of a number of systematic errors thanks to dedicated mock data. This results in a roughly $\sim 10\%$ stronger constraining power on $\Lambda$CDM parameters, and importantly, the lensing anomaly seems to be significantly reduced or even absent in those data. In fact, it has been shown that they lead to weaker bounds to $\sum m_{\nu}$ than the Plik 2018 likelihood, see Rosenberg et al. (2022); Tristram et al. (2024); Allali and Notari (2024).
@@
Figure 2: Implications of BAO measurements of $\Omega_{m}$ and $H_{0}r_{d}$ for $\sum m_{\nu}$ inferences. We show the posterior density contours using Planck data (grey and dots), as well as the regions favoured by the full SDSS BAO sample (in blue), DESI-Y1 (in red), and DESI-Y1 without the $z=0.7$ outliers (in black dashed). We can clearly see that the outliers at $z=0.7$ have a relevant shift on the parameter space with implications for the neutrino mass.
@@ III.1 Cosmological Data: CMB, BAO and Supernova
In what follows, we will first perform a comprehensive analysis of the bound on neutrino masses coming from considering Planck data alone, to highlight the role of potential anomalies (whether a statistical fluke or a systematic effect) in the data, and how subsequent data release have affected those bounds. To that end, we consider the following likelihood combinations:
@@
  • DESI-Y1-no07 – We consider the full set of BAO data as reported in Adame et al. (2024b, a) but without including the two data points at $z=0.7$. Since these data points are in $\sim\!3\sigma$ tension [^5] Out of the 22 data points in the DESI-Y1 BAO sample, most of them are in good agreement with Planck $\Lambda$CDM predictions, but these $z=0.7$ outliers feature $\sim 2.6\sigma$ and $\sim 2.8\sigma$ deviations. The chances of this occurring (assuming Gaussian and uncorrelated errors) is of only $\sim 1.8\%$, see Efstathiou (2024). with Planck predictions, it is interesting to explore their impact on the $\sum m_{\nu}$ bound to understand what would happen if these outliers are not confirmed by future data, as highlighted in Figure 2.
@@ III.2 Analysis methodology: comparing Bayesian and frequentist framework
As an alternative to the Bayesian approach, and to compare and contrast the results, we make use of the frequentist framework, relying on a series of $\chi^{2}$ optimization at fixed neutrino mass, to build a profile likelihood and derive confidence intervals. The main advantages of the frequentist approach are that i) constraints are insensitive to the specific choice of priors, and ii) the presence of the physical boundary can be accounted for in a statistically consistent manner Feldman and Cousins (1998). Thus, a direct comparison between the two approaches may provide a better handle on the relevance of these subtleties, as well as on the robustness of the cosmological bound.
@@
On the Bayesian side, we perform a Markov Chain Monte Carlo (MCMC) sampling of the posterior distribution, using the publicly available MontePython code Brinckmann and Lesgourgues (2018); Audren et al. (2013). Cosmological and nuisance parameters are varied according to the “fast” and “slow” parameters decomposition Lewis (2013). The chains are then marginalized with GetDist Lewis (2019) in order to extract the bounds. On the frequentist side, instead of posterior marginalization, the relevant procedure is likelihood profiling, which requires the minimization of the $\chi^{2}$ function for a fixed value of the parameter of interest ($\sum m_{\nu}$ in our case), varying simultaneously the $N-1$ remaining cosmological and nuisance parameters. In particular, the various likelihood combinations described in Section III contain of the order of $N\sim 20-25$ parameters making the minimization a highly non-trivial task. In this context, we have opted to perform the numerical minimization with the simulated annealing algorithm implemented in Procoli Karwal et al. (2024) [^6] We note that there are other public codes available such as Prospect Holm et al. (2023) and CAMEL Henrot-Versillé et al. (2016). While Prospect also relies on a simulated annealing and would likely provide similar results as Procoli, CAMEL relies on the quasi-Newtonian optimizer Minuit James and Roos (1975). We have not been able to obtain converged likelihood profiles with CAMEL, highlighting one of the difficulty in the frequentist approach.. Although this procedure is straightforward, the HiLLiPoP likelihood requires significantly more care to successfully converge. We attribute this to the fact that HiLLiPoP is the sole CMB likelihood that is not binned, and is therefore noisier than the others. For this reason, our profiling is not performed over a regular grid of $\sum m_{\nu}$. This is, however, not a major cause for concerns, since to properly calibrate the frequentist test statistics we will in any case fit the likelihood profiles. In addition, we can extrapolate the resulting fits to the unphysical negative neutrino mass region, allowing us to asses the potential preference for negative masses recently displayed by some datasets and discussed in Craig et al. (2024); Green and Meyers (2024); Elbers et al. (2024).
@@ IV.1 Planck Only Analyses
Figure 3: Profile likelihoods for the neutrino masses within $\Lambda$CDM for three different versions of the Planck likelihoods: Planck18-PR3, CamSpec22-PR4, and HiLLiPoP23-PR4. For comparison purposes we also show the Planck 2013 results from Ade et al. (2014) in red where the potential trend for a best fit in the “negative” regime was first highlighted. We clearly see that the bound on the neutrino masses changes significantly for each implementation of the likelihood, being HiLLiPoP the one giving the looser constraints.
Planck only (no lensing) $95\%$ C.L. $\sum m_{\nu}$ (eV)
@@
Table 1: Upper limits at 95% CL on the neutrino mass within $\Lambda$CDM using various Planck likelihoods (no lensing likelihood included). We show the Bayesian limits compared with the two frequentist approaches (B.L. = Bounded Likelihood, and F.C. = Feldman-Cousins) described in Section III.2. We do not quote frequentist F.C. bounds for Planck18-PR3 and CamSpec22-PR4 due to their non-Gaussian behaviour in the physical region, which precludes a reliable extrapolation into the unphysical region.
Figure 4: Neutrino mass profile likelihoods using the full Planck temperature and polarization data $\Lambda$CDM allowing to vary the unphysical $A_{\rm lens}$ parameter which is strongly correlated with $\sum m_{\nu}$. We can see that the bounds are significantly relaxed and comparable to the KATRIN laboratory upper limit.
@@ IV.2 Planck + BAO:
In particular, as stated in the introduction, one of the main goals of our study is to understand why the bound from Planck+DESI-Y1 is so strong, and whether it really points to negative neutrino masses. To understand this, in Fig. 5 we show the analyses of Planck data [^7] We stress that Lensing-PR3 is now included in the analysis. combined with: (a) DESI-Y1 BAO results, (b) the full SDSS combination (which have a similar statistical power in the $\Omega_{m}$ vs $H_{0}r_{d}$ plane as DESI-Y1, see Fig. 2), (c) the DESI/SDSS BAO combination that takes the BAO results at each redshift bin from the survey that has greatest statistical power, and (d) the DESI-Y1 BAO combination but without the two outliers at $z=0.7$ which deviate by $\sim 3\sigma$ from $\Lambda$CDM predictions. Importantly, in all data set combinations, we consider separately the three latest Planck likelihood implementations.
@@
First, from panel (a), we confirm that the extrapolation of the results from Planck+DESI-Y1 to the unphysical region seem to favor negative masses. However, in our approach, given uncertainties in the profile represented by the bands, the preference remains at low statistical significance. In fact, we find that it is the combination with HiLLiPoP23-PR4 that seems to lead to the largest preference (reaching at most $\sim 1\sigma$). This may appear counter-intuitive given results presented in previous section. We attribute this to the fact that HiLLiPoP23-PR4 has the weakest constraining power on neutrino masses, and thus the statistical power from DESI can be more pronounced. Second, by comparing the results from Planck+DESI-Y1 and Planck+SDSS in panel (b), one can clearly see that the bound on the neutrino mass are significantly weaker in the latter case and that the preference for negative neutrino masses further decrease. This is particularly striking for HiLLiPoP23-PR4 that is now fully compatible with positive masses. However, when comparing the Planck+DESI-Y1 and Planck+DESI-Y1+SDSS (panel (c)) one notices that the profiles are quite similar and only slightly shifted towards more positive values for the latter case. This suggests that the low-$z$ data points from DESI, that are replaced by SDSS is this analysis, do not play a significant role in the preference for negative neutrino masses. Rather, we can compare the results from Planck+DESI-Y1 and Planck+DESI-Y1no07 (panel (d)) which excludes the outliers at $z=0.7$. We find that removing these data points loosens the bound and also makes all likelihood combinations to peak at around $\sum m_{\nu}\simeq 0$, thus removing the potential preference for negative neutrino masses. Although it is obvious that removing two data points from DESI-Y1 is not a valid statistical procedure, this exercise shows that these outliers have a very significant impact on the neutrino mass bound. We note that the DESI-Y1+SDSS BAO combination does include these outliers in the combination as the effective volume of DESI is already larger than that of SDSS at $z=0.7$.

### v2

@@ Abstract
Cosmological neutrino mass bounds are becoming increasingly stringent. The latest limit within $\Lambda$CDM from Planck 2018+ACT lensing+DESI is $\sum m_{\nu}<0.072\,{\rm eV}$ at 95% CL, very close to the minimum possible sum of neutrino masses ($\sum m_{\nu}>0.06\,{\rm eV}$), hinting at vanishing or even “negative” cosmological neutrino masses. In this context, it is urgent to carefully evaluate the origin of these cosmological constraints. In this paper, we investigate the robustness of these results in three ways: i) we check the role of potential anomalies in Planck CMB and DESI BAO data; ii) we compare the results for frequentist and Bayesian techniques, as very close to physical boundaries subtleties in the derivation and interpretation of constraints can arise; iii) we investigate how deviations from $\Lambda$CDM, potentially alleviating these anomalies, can alter the constraints. From a profile likelihood analysis, we derive constraints in agreement at the $\sim 10\%$ level with Bayesian posteriors. We find that the weak preference for negative neutrino masses is mostly present for Planck 18 data, affected by the well-known ‘lensing anomaly’. It disappears when the new Planck 2020 HiLLiPoP is used, leading to significantly weaker constraints. Additionally, the pull towards negative masses in DESI data stems from the $z=0.7$ bin, which contains a BAO measurement in $\sim 3\sigma$ tension with Planck expectations. Without this bin, and in combination with HiLLiPoP, the bound relaxes to $\sum m_{\nu}<0.11\,{\rm eV}$ at 95% CL. The recent preference for dynamical dark energy alleviates this tension and further weakens the bound. As we are at the dawn of a neutrino mass discovery from cosmology, it will be very exciting to see if this trend is confirmed by future data.
@@ I Introduction: Neutrino Mass bounds as of mid 2024
Clearly, the current cosmological limit is very close to the minimal value in normal ordering and already disfavours to some degree the inverted one. Importantly, the limit in Eq. (1) is so close to the minimum physical boundary that statistical statements about the neutrino mass need to be taken with care. However, cosmological limits are typically derived within a Bayesian framework, and as such will depend upon the priors used. While the dependence may be weak when the likelihood largely dominates over the prior, and far away from physical boundaries, it has already been established that neutrino masses are strongly sensitive to the choice of prior Simpson et al. (2017); Gariazzo et al. (2018, 2022, 2023), even with DESI Adame et al. (2024a). Moreover, and perhaps even more surprising, there is still no hint of a non-zero mass in the posterior probability density, and in fact, it has been argued that cosmological data may favor a negative effective neutrino mass Craig et al. (2024); Green and Meyers (2024); Elbers et al. (2024). While this could be the result of a statistical fluctuation or a systematic effect, it could potentially be the indication of new phenomena in cosmology with groundbreaking implications.
@@
The rest of our study is structured as follows: first, in Section II we briefly review the main cosmological implications of massive neutrinos. We will highlight the crucial role of Planck data as well as why BAO data can significantly tighten the constraints on the neutrino mass. In Section III we outline the various data sets that we will use, and describe the statistical procedure (both frequentist and Bayesian) that we use to analyse the data and derive bounds. In Section IV we present our main results focusing first on analyses with Planck data only, exploring in detail the role of the lensing anomaly; we then perform comprehensive analysis of Planck and BAO data, comparing in particular DESI and SDSS; and finally, check the impact of considering extensions to $\Lambda$CDM. In Section V we specifically address the potential cosmological preference for a negative neutrino mass. Finally, in Section VI we draw our conclusions. Additional material supporting our findings, including posterior distributions, correlations and comparisons with previous works, is provided in the Appendices.
@@ II Cosmological Impact of the neutrino mass
[…]
@@
Importantly, the CamSpec and HiLLiPoP likelihoods have been recently updated since the 2018 analysis, in light of a new set of maps produced by the Planck collaboration called NPIPE. The NPIPE maps exploit a number of improvements in the processing of time ordered data to allow for an increase in the signal-to-noise ratio at small scales. They also allow to use a larger sky fraction, and incorporate a better handling of a number of systematic errors thanks to dedicated mock data. This results in a roughly $\sim 10\%$ stronger constraining power on $\Lambda$CDM parameters, and importantly, the lensing anomaly seems to be significantly reduced or even absent in those data. In fact, it has been shown that they lead to weaker bounds to $\sum m_{\nu}$ than the Plik 2018 likelihood, see Rosenberg et al. (2022); Tristram et al. (2024); Allali and Notari (2024).
@@
Figure 2: Implications of BAO measurements of $\Omega_{m}$ and $H_{0}r_{d}$ for $\sum m_{\nu}$ inferences. We show the posterior density contours using Planck data (grey and dots), as well as the regions favoured by the full SDSS BAO sample (in blue), DESI-Y1 (in red), and DESI-Y1 without the $z=0.7$ bin, which contains a $2.6\sigma$ outlier (in black dashed). We can clearly see that the $z=0.7$ leads to a relevant shift on the parameter space with implications for the neutrino mass.
@@ III.1 Cosmological Data: CMB, BAO and Supernova
In what follows, we will first perform a comprehensive analysis of the bound on neutrino masses coming from considering Planck data alone, to highlight the role of potential anomalies (whether a statistical fluke or a systematic effect) in the data, and how subsequent data releases have affected those bounds. To that end, we consider the following likelihood combinations:
@@
  • DESI-Y1-no07 – We consider the full set of BAO data as reported in Adame et al. (2024b, a) but without including the two data points at $z=0.7$. Since this corresponds precisely to the redshift of dark energy domination and this bin contains a data point in $\sim\!3\sigma$ tension [^5] Out of the 22 BAO data points, most of them are in good agreement with Planck $\Lambda$CDM predictions, but there are two DESI-Y1 data points at $z=0.5$ and $z=0.7$ which deviate by $\sim 2.8\sigma$ and $\sim 2.6\sigma$, respectively. Given a total of 22 total measurements of BAO, the chances of this occurring (assuming Gaussian and uncorrelated errors) is of only $\sim 1.8\%$, see Efstathiou (2024). We further note that the DESI-Y1 BAO measurements at $z=0.7$ are also in some tension with the old SDSS ones at similar redshifts. with Planck predictions, it is interesting to explore its impact on the $\sum m_{\nu}$ bound to understand what would happen if these outliers are not confirmed by future data, as highlighted in Figure 2.
@@ III.2 Analysis methodology: comparing Bayesian and frequentist framework
As an alternative to the Bayesian approach, and to compare and contrast the results, we make use of the frequentist framework, relying on a series of $\chi^{2}$ optimizations at fixed neutrino mass, to build a profile likelihood and derive confidence intervals. The main advantages of the frequentist approach are that i) constraints are insensitive to the specific choice of priors, and ii) the presence of the physical boundary can be accounted for in a statistically consistent manner Feldman and Cousins (1998). Thus, a direct comparison between the two approaches may provide a better handle on the relevance of these subtleties, as well as on the robustness of the cosmological bound.
@@
On the Bayesian side, we perform a Markov Chain Monte Carlo (MCMC) sampling of the posterior distribution, using the publicly available MontePython code Brinckmann and Lesgourgues (2018); Audren et al. (2013). Cosmological and nuisance parameters are varied according to the “fast” and “slow” parameters decomposition Lewis (2013). The chains are then marginalized with GetDist Lewis (2019) in order to extract the bounds. On the frequentist side, instead of posterior marginalization, the relevant procedure is likelihood profiling, which requires the minimization of the $\chi^{2}$ function for a fixed value of the parameter of interest ($\sum m_{\nu}$ in our case), varying simultaneously the $N-1$ remaining cosmological and nuisance parameters. In particular, the various likelihood combinations described in Section III contain of the order of $N\sim 20-25$ parameters making the minimization a highly non-trivial task. In this context, we have opted to perform the numerical minimization with the simulated annealing algorithm implemented in Procoli Karwal et al. (2024) [^6] We note that there are other public codes available such as Prospect Holm et al. (2023) and CAMEL Henrot-Versillé et al. (2016). While Prospect also relies on a simulated annealing algorithm and would likely provide similar results as Procoli, CAMEL relies on the quasi-Newtonian optimizer Minuit James and Roos (1975). We have not been able to obtain converged likelihood profiles with CAMEL, highlighting one of the difficulties in the frequentist approach.. Although this procedure is straightforward, the HiLLiPoP likelihood requires significantly more care to successfully converge. We attribute this to the fact that HiLLiPoP is the sole CMB likelihood that is not binned, and is therefore noisier than the others. For this reason, our profiling is not performed over a regular grid of $\sum m_{\nu}$. This is, however, not a major cause for concerns, since to properly calibrate the frequentist test statistics we will in any case fit the likelihood profiles. In addition, we can extrapolate the resulting fits to the unphysical negative neutrino mass region, allowing us to asses the potential preference for negative masses recently displayed by some datasets and discussed in Craig et al. (2024); Green and Meyers (2024); Elbers et al. (2024).
@@ IV Cosmological neutrino mass bounds: A frequentist vs Bayesian comparison
Figure 3: Profile likelihoods for the neutrino masses within $\Lambda$CDM for three different versions of the Planck likelihoods: Planck18-PR3, CamSpec22-PR4, and HiLLiPoP23-PR4. For comparison purposes we also show the Planck 2013 results from Ade et al. (2014) in red where the potential trend for a best fit in the “negative” regime was first highlighted. We clearly see that the bound on the neutrino masses changes significantly for each implementation of the likelihood, being HiLLiPoP the one giving the looser constraints. Solid and dashed lines correspond to parabolic fits where the $\Delta\chi^{2}$ points up to 4 or 2 were used in the fit, respectively.
@@ IV.1 Planck Only Analyses
In Fig. 5 we show the results but now including the Lensing-PR3 lensing likelihood and with the physical condition $A_{\rm lens}=1$. By comparing the results with those without lensing in Fig. 3 we can clearly see that while the impact on the Planck18-PR3 analysis is mild at the level of the $2\sigma$ limit, the extrapolated behaviour to negative neutrino masses shows significantly weaker support for a negative best fit. This is because the lensing likelihood does not feature a lensing anomaly. The shift for the CamSpec22-PR4 and HiLLiPoP23-PR4 cases is substantially more pronounced and as shown in Table 1 the 95% CL limit improves by a factor of 1.5 when the lensing likelihood is added. We note that for the data set combination HiLLiPoP23-PR4+Lensing-PR3 the minimization procedure was rather challenging highlighting that a non-global minimum of the $\chi^{2}$ at around $\sum m_{\nu}\sim 0.2\,{\rm eV}$ may be present. In Appendix A we show the Bayesian posterior in Fig. 11 and there is a priori no evidence for multi-modality in it.
Figure 4: Neutrino mass profile likelihoods using the full Planck temperature and polarization data for $\Lambda$CDM allowing to vary the unphysical $A_{\rm lens}$ parameter which is strongly correlated with $\sum m_{\nu}$. We can see that the bounds are significantly relaxed and comparable to the KATRIN laboratory upper limit.
Figure 5: Neutrino mass profile likelihoods using the full Planck temperature, polarization and lensing data for $\Lambda$CDM. This should be compared with Fig. 3 that does not include lensing.
Planck only $95\%$ C.L. $\sum m_{\nu}$ (eV)
@@
With lensing likelihood
$\Lambda$CDM Planck18-PR3 0.25 0.20 0.18
CamSpec22-PR4 0.21 0.19 0.18
HiLLiPoP23-PR4 0.34 0.33 0.32
Table 1: Upper limits at 95% CL on the neutrino mass within $\Lambda$CDM using various Planck likelihoods. We show the Bayesian limits compared with the two frequentist approaches (B.L. = Bounded Likelihood, and F.C. = Feldman-Cousins) described in Section III.2. We do not quote frequentist F.C. bounds for Planck18-PR3 and CamSpec22-PR4 (no lensing) due to their non-Gaussian behaviour in the physical region, which precludes a reliable extrapolation into the unphysical regime.
@@ IV.2 Planck + BAO:
Figure 6: Reconstructed posterior distribution of $\sum m_{\nu}$ in a mock analysis of Planck18-PR3+Lensing+BAO with three configurations: $\sum m_{\nu}=0,0.06,0.1$ eV. We also show the posterior reconstructed from the real data for comparison.
@@
In particular, as stated in the introduction, one of the main goals of our study is to understand why the bound from Planck+DESI-Y1 is so strong, and whether it really points to negative neutrino masses.
In order to gauge how unexpected is the result from Planck+DESI-Y1, it is instructive to first estimate the sensitivity of this data combination to neutrino mass, as for instance done in Ref. Brinckmann et al. (2019). To do so, we perform a mock data analysis, making use of the MontePython and the Fake__\__planck__\__realistic likelihood Brinckmann et al. (2019), that we combine with our own mock likelihood of DESI-Y1, that simply makes use of the covariance matrix from the data, replacing the measurements with a fiducial model prediction. For the fiducial, we study three configurations: $\sum m_{\nu}=0,0.06,0.1$ eV, fixing the other $\Lambda$CDM parameters as reconstructed from Planck18-PR3+Lensing+BAO Aghanim et al. (2020a). We run MCMC chains, and show the reconstructed posteriors in Fig. 6. It is clear that the real data (black line) provide significantly stronger constraints than expected, and in fact, our mock data suggests that the combination of Planck+DESI-Y1 has the sensitivity necessary to detect at small significance non-zero neutrino masses, assuming these respect the laboratory lower limits. Barring issue with our mock likelihoods, it is thus clear that there is some ‘anomaly’ in the real data (whether due to a fluke, or new physics). As already discussed and as we will elaborate in the following with analyses of the real data, the two results that seem to be driving this unexpected constraints beyond the sensitivity are Planck likelihoods through the lensing anomaly and the DESI-Y1 data at $z=0.7$. Indeed, when we compare instead the posterior for HiLLiPoP, with a significantly reduced lensing anomaly, and remove the DESI outliers, the posterior is much closer to the expected sensitivity as shown by the purple line in Fig. 6. When in presence of such anomalous behaviour from the actual data when compared to the expected sensitivities, an alternative method to the Feldman-Cousins analysis was proposed by Lokhov and Tkachov Lokhov and Tkachov (2015) ensuring also proper coverage when deriving the confidence intervals but conservatively replacing the anomalous results pushing into the unphysical region with the expected sensitivity in absence of a signal. Since we find that the main contributors driving the anomalous results can be identified and that their impact can be almost entirely removed for instance through the HiLLiPoP + DESI-Y1 combination, we prefer not to implement this method as it would hide the different behaviours of the different datasets and cosmological models we want to analize and compare here. However, we do find the comparison with the expected sensitivities presented in Fig. 6, on which the LT method is based, very illustrating. Furthermore, we can compare our expected sensitivities of Planck+DESI-Y1 to the ones obtained in previous forecasts using the full DESI reach Brinckmann et al. (2019). We find a $1\sigma$ sensitivity of $\sim\!0.06\,{\rm eV}$ while for Planck+DESI-full Ref. Brinckmann et al. (2019) reports a $1\sigma$ sensitivity of $0.04\,{\rm eV}$. This highlights what is the level of improvement expected from the additional 5 years of BAO data, although an even higher sensitivity can of course be obtained from a full shape analysis of the matter power spectrum, see Font-Ribera et al. (2014); Amendola et al. (2013); Aghamousa et al. (2016); Chudaykin and Ivanov (2019).
@@
We now turn to the real data. In Fig. 7 we show the analyses of Planck data [^7] We stress that Lensing-PR3 is now included in the analysis. combined with: (a) DESI-Y1 BAO results, (b) the full SDSS combination (which have a similar statistical power in the $\Omega_{m}$ vs $H_{0}r_{d}$ plane as DESI-Y1, see Fig. 2), (c) the DESI/SDSS BAO combination that takes the BAO results at each redshift bin from the survey that has greatest statistical power, and (d) the DESI-Y1 BAO combination but without the $z=0.7$ bin that contains data which deviate by $\sim 3\sigma$ from $\Lambda$CDM predictions. Importantly, in all data set combinations, we consider separately the three latest Planck likelihood implementations.
@@
First, from panel (a), we confirm that the extrapolation of the results from Planck+DESI-Y1 to the unphysical region seem to favor negative masses. However, in our approach, given uncertainties in the profile represented by the bands, the preference remains at low statistical significance. In fact, we find that it is the combination with HiLLiPoP23-PR4 that seems to lead to the largest preference (reaching at most $\sim 1\sigma$). This may appear counter-intuitive given results presented in previous section. We attribute this to the fact that HiLLiPoP23-PR4 has the weakest constraining power on neutrino masses, and thus the statistical power from DESI can be more pronounced. Second, by comparing the results from Planck+DESI-Y1 and Planck+SDSS in panel (b), one can clearly see that the bound on the neutrino mass are significantly weaker in the latter case and that the preference for negative neutrino masses further decrease. This is particularly striking for HiLLiPoP23-PR4 that is now fully compatible with positive masses. However, when comparing the Planck+DESI-Y1 and Planck+DESI-Y1+SDSS (panel (c)) one notices that the profiles are quite similar and only slightly shifted towards more positive values for the latter case. This suggests that the low-$z$ data points from DESI, that are replaced by SDSS is this analysis, do not play a significant role in the preference for negative neutrino masses. Rather, we can compare the results from Planck+DESI-Y1 and Planck+DESI-Y1no07 (panel (d)) which excludes the outliers at $z=0.7$. We find that removing these data points loosens the bound and also makes all likelihood combinations to peak at around $\sum m_{\nu}\simeq 0$, thus removing the potential preference for negative neutrino masses. Although it is obvious that removing two data points from DESI-Y1 is not a valid statistical procedure, this exercise shows that DESI-Y1 BAO data at $z=0.7$ have a very significant impact on the neutrino mass bound. We note that the DESI-Y1+SDSS BAO combination does include these outliers in the combination as the effective volume of DESI is already larger than that of SDSS at $z=0.7$.